# StrandHead: Text to Strand-Disentangled 3D Head Avatars Using Hair Geometric Priors

### v1

@@ 3.4 Strand-Based Hair Generation
where $\hat{p}^{i}_{j}$ and ${p}^{i}_{j}$ are the position of the $j$-th point on the $i$-th polyline of the GT and generated hair, respectively, and $\hat{o}^{i}_{j}$, $o^{i}_{j}$, $\hat{c}^{i}_{j}$, and $c^{i}_{j}$ denote their orientation and curvature, respectively. Here, orientation $o^{i}_{j}=(p^{i}_{j+1}-p^{i}_{j})/\|p^{i}_{j+1}-p^{i}_{j}\|_{2}$ represents the direction of change in strand position, and curvature $c^{i}_{j}=\|o^{i}_{j}-o^{i}_{j-1}\|_{2}$ denote the rate of change in strand orientation. Given the hair description, we select an optimal pre-trained NST as initialization and further optimize it in the following stages.
@@
Differentiable Prismatization. Actually, as there is no reliable strand-based differentiable renderer, performing end-to-end hair strands optimization becomes challenging. Further, since we have no constrained multi-view images, we cannot directly model the hair geometry and appearance using 3DGS [20] or line-based renderer [58], as existing hair reconstruction methods [58, 69, 33, 88]. A feasible alternative is to first convert the strands into quad meshes [87], similar to NeuralHaircut [66], and then leverage mesh optimization frameworks to model hair strands. However, this non-watertight stripe-like mesh easily produces ambiguous normal maps or excessively thin sides, which significantly reduce the stability of optimization (Sec. 4.4). Therefore, a more stable strand-to-mesh conversion method with superior optimization properties is urgently necessary.
@@
To this end, we propose a novel differentiable prismatization algorithm inspired by the internal structure of hair (i.e., hair fiber is a dielectric cylinder covered with tilted scales and with a pigmented interior [19, 35]). This algorithm can efficiently and differentiably convert strands into watertight prismatic meshes with arbitrary thickness and lateral edges to approximate the cylindrical structure of hair, enabling the use of mesh-based renderers and algorithms for strand-level modeling. Specifically, given a hair strand $s$, our differentiable prismization algorithm converts it into a watertight prismatic mesh with $K$ lateral edges and radius $R$ through the following five steps:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
Figure 3 shows an example of converting a hair strand into an octagonal prism mesh. Importantly, the proposed differentiable prismatization algorithm can be easily implemented on GPU, achieving flexible and fast prismatization of hair strands and paving a new way for hair modeling. The algorithm details are shown in the Supp. Mat.
Figure 3: Illustration of the process of converting a hair strand into an octagonal prism mesh using the differentiable prismatization algorithm.
Hair Geometry and Texture Modeling. With the help of our proposed differentiable prismatization algorithm, we create the hair geometry and texture in two stages as Sec. 3.3 using the following losses:
@@
$\nabla_{T}\mathcal{L}^{\text{hn}}_{SDS}$ $=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{\text{hn}}}(n^{\text{h+s}}_{t};y_{\text{h}},t)-\epsilon)\frac{\partial n^{\text{h+s}}}{\partial T}\right],$ (9)
@@
$\nabla_{T}\mathcal{L}^{\text{hd}}_{SDS}$ $=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{\text{hd}}}(d^{\text{h+s}}_{t};y_{\text{h}},t)-\epsilon)\frac{\partial d^{\text{h+s}}}{\partial T}\right],$ (10)
$\nabla_{{\psi}_{\text{s}}}\mathcal{L}^{\text{hc}}_{SDS}$ $=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{\text{hc}}}(c^{\text{h+s}}_{t};n^{\text{h+s}},y_{\text{h}},t)-\epsilon)\frac{\partial c^{\text{h+s}}}{\partial{\psi}_{\text{s}}}\right],$ (11)
$\nabla_{{\psi}_{\text{s}}}\mathcal{L}^{\text{hc}}_{MSDS}$ $=\mathbb{E}_{t,\epsilon}\left[(H(c^{\text{h+s}}_{t};n^{\text{h+s}},y_{\text{h}},t)-\epsilon)\frac{\partial c^{\text{h+s}}}{\partial{\psi}_{\text{s}}}\right],\vspace{-1cm}$ (12)
@@
where $n^{\text{h+s}}$, $d^{\text{h+s}}$, and $c^{\text{h+s}}$ are the rendered normal, depth, and color maps of the head $\theta_{\text{h}}$ with strand-based hair prismatic mesh $M_{p}=DP(G(T))$, respectively. Here $DP(\cdot)$ denotes the differentiable prismatization operation function, ${\psi}_{\text{s}}$ is the strand texture field.
Figure 4: The distribution of $O_{\text{ori}}$ and $C_{\text{mean}}$ in the USC-HairSalon dataset [15]. The results indicate that (1) neighboring strand orientations are highly consistent; (2) strand curvature is strongly and positively related to the haircut curliness.
@@
Prior-Driven Losses. However, due to the excessive flexibility of hair strands, relying solely on the SDS loss leads to unnatural strand orientations and unrealistic hairstyles (Sec. 4.4). Further, the lack of constrained multi-view images prevents us from leveraging hair-specific features [45] to provide powerful supervision as in previous works [58].
Figure 5: Examples of high-fidelity and diverse 3D heads and strand-accurate haircuts generated by our method. The upper visualization includes rendered color and normal maps of the head and hair prism meshes. The lower visualization shows the physics-based rendering result using Blender [10]. Please zoom in for detailed views, and refer to the Supp. Mat. for video demonstrations.
@@
To address this issue, we propose two straightforward yet solid prior-driven losses based on observations of the hair strand orientation and curvature to ensure the rationality of the generated hair strands. Specifically, we observe two geometric properties of hair: (1) neighboring strand orientations are highly consistent, and (2) strand curvature is strongly and positively correlated with the curliness of the hairstyle. To validate the properties, we calculate the cosine similarity of adjacent hair strand orientations $CS_{\text{ori}}$ and the average curvature $C_{\text{mean}}$ of 343 hairstyles in the USC-HairSalon dataset [15] using the following equations and visualize their distributions and some examples in Fig. 4:
@@
$C_{\text{mean}}$ $=\frac{1}{N_{s}N_{p}}\sum^{N_{s}}_{i=1}\sum^{N_{p}}_{j=1}c^{i}_{j},$ (14)
where $A(i)$ is the set of adjacent strands for the $i$-th strand, and $|A(i)|$ is the number of neighboring strands.
@@
$\mathcal{L}_{\text{cur}}$ $=\|C_{\text{mean}}-C_{\text{target}}\|_{1},$ (16)
where $C_{\text{target}}$ represents the target average curvature set according to the input hair description. These two losses ensure the rationality of generated strands by supervising the consistency of orientations between adjacent strands and the overall hairstyle’s curvature.
Additionally, we introduce a series of geometry-aware losses $\mathcal{L}_{bbox}$, $\mathcal{L}_{face}$, and $\mathcal{L}_{colli}$, which prevent the hair from exceeding the bounding box, obscuring the face, and colliding with the head, respectively. Details are provided in the Supp. Mat. The final loss function for optimizing hair geometry is expressed as follows:
@@
Strand-Aware Texture Field. Since the complexity of hair textures, a basic texture field cannot fully capture the lifelike appearance of strands (Sec. 4.4). Therefore, we propose a strand-aware texture field that better models high-frequency color variations. Specifically, we introduce two improvements to the basic texture field: first, we replace Euclidean coordinates with scalp UV coordinates which are more uniformly distributed. Second, we incorporate strand orientations as additional input information to model orientation-dependent texture. These enhancements enhance the realism of the generated results by better creating the high-frequency appearance variations, but also ensure consistent colors across different faces of a single prismatic mesh since the input features are strand-based. For more details, please refer to the Supp. Mat.
@@ 3.5 Implementation Details
Our approach utilizes PyTorch [47] and ThreeStudio [11] for implementation. Experiments are carried out on an Ubuntu server with A6000 GPUs. Starting with a textual description of a head with hair, such as “A man with brown hair”. we initially create a bald head using the prompt “A bald man” and then generate hair based on “A man with brown hair”. Further implementation details are available in the Supp. Mat.
@@
Figure 7: Qualitative comparison with HAAR [67]. For better visual comparison, we interpolate the hairstyles to approximately 10,000 strands and apply a consistent appearance. Since HAAR does not model heads, its generated results frequently display hair-head collisions, highlighted within the black box.
@@ 4 Experiments
Some examples of 3D head models with strand-based hair generated by StrandHead are shown in Fig. 5. Due to space limitations, only the key experimental settings and results are presented here. For more results and details, please refer to our Supp. Mat.
@@ 4.1 Experimental Settings
Baselines. We compare StrandHead with the SOTA methods for head avatar and haircut generation. The baseline methods for head generation include text-to-holistic-head works (HeadArtist [30], HumanNorm [17], and HeadStudio [99]) and text-to-decoupled-head work (TECA [89]). For text-to-hair generation, we consider general text-to-3D methods (MVDream [65], GaussianDreamer [85], LucidDreamer [28], and RichDreamer [51]), strand-specific text-to-hair methods (HAAR [67]), and hair-disentangled text-to-head methods (TECA [89]).
@@
Evaluation Metrics. Current text-to-3D methods typically use CLIP-based metrics to assess text-image alignment and the quality of generated outputs. However, research has shown [16, 32] that CLIP-based metrics fall short in capturing the fine-grained alignment between 3D content and input prompts, as further validated by our experiments in the Supp. Mat. To address this limitation, and drawing inspiration from Progressive3D [7] and Barbie [68], we use BLIP-VQA [24, 23] and BLIP2-VQA [25, 23] to evaluate both current methods and StrandHead’s generative capabilities. Specifically, we convert each prompt into two questions to separately verify head and hair. We then feed the rendered image of the generated 3D content into the VQA model, asking each question in sequence and using the probability of a “yes” response as our evaluation metric. For example, the head prompt “A man with black hair” is transformed into “Is the person in the picture a man?” and “Does the person in the picture have black hair?” The hair prompt “Black hair” becomes “Is the object in the picture black hair?” Finally, we conduct a user study by randomly selecting 10 generated examples and asking 30 volunteers to assess (1) generation quality and (2) text-image alignment, and select the best methods.
@@ 4.2 Comparisons of Head Generation
As illustrated in the upper section of Tab. 2, our approach outperforms all comparison methods across every evaluation metric. The qualitative results on the left side of Fig. 6 further highlight the advantages of our method. Compared to other head generation methods, the head avatars generated by StrandHead exhibit not only more refined facial geometry and texture details but separable strand-accurate hair with a plausible appearance that integrates seamlessly with physics-based rendering and simulation systems. To the best of our knowledge, StrandHead is the first head generation framework to achieve strand-level hair modeling, a capability that holds substantial potential for advancing human-centric 3D AIGC applications in the industry.
@@ 4.3 Comparisons of Hair Generation
Unlike the text-to-strand approach HAAR [67], our approach does not require 3D hair training data, enabling it to better generate uncommon hairstyles (e.g., slicked-back and side-swept hairstyles), as further validated by the physics-based rendering comparison in Fig. 7. Additionally, StrandHead can generate both hair textures and head models, which not only extends the range of downstream applications but also avoids hair-head collisions (see Fig. 7).
Figure 8: Ablation study on (a) strand-level optimization, (b) orientation consistency loss, (c) curvature regularization loss, (d) differentiable prismatization, and (e) strand-aware texture field.
@@ 4.4 Ablation Study
Effect of Strand-Level Optimization. As illustrated in Fig. 8-(a), under dual constraints of SDS loss and prior-driven losses, the optimized hair better aligns with the text and integrates harmoniously with the human head.
Effect of Orientation Consistency Loss. Fig. 8-(b) demonstrates that incorporating $\mathcal{L}_{\text{cur}}$ provides strong guidance for generating realistic and natural hair.
Effect of Curvature Regularization Loss. As shown in Fig. 8-(c), our $\mathcal{L}_{\text{cur}}$ preserves the desired curliness by constraining the strand curvature.
Effect of Differentiable Prismatization. As shown in Fig. 8-(d), compared to the quad mesh used by NeuralHaircut [66] our proposed differentiable prismatization algorithm enables more stable gradient backpropagation for reliable strand-based hair optimization, while also avoiding abnormal normal issues caused by non-watertight meshes.
Effect of Strand-Aware Texture Field. As shown in Fig. 8-(e), our proposed strand-aware texture field accurately models high-frequency appearance details by switching coordinate spaces and incorporating orientation information, resulting in more realistic strand textures.
Figure 9: Applications of StrandHead.
4.5 Applications
As illustrated in Fig. 9-(a), StrandHead enables the flexible combination of any head with any hairstyle, greatly enhancing the playability and reusability of head generation. Furthermore, due to our precise strand-level hair modeling, we can not only edit hairstyles by manipulating strand geometry and texture but also leverage mature industrial software to achieve physics-based realistic rendering and simulation, as shown in Fig. 9-(b) and (c).

### v3

@@ 3.4 Hair Geometry Generation
where $\hat{p}^{i}_{j}$ and ${p}^{i}_{j}$ are the position of the $j$-th point on the $i$-th polyline of the GT and generated hair, respectively, and $\hat{o}^{i}_{j}$, $o^{i}_{j}$, $\hat{c}^{i}_{j}$, and $c^{i}_{j}$ denote their orientation and curvature, respectively. Here, orientation $o^{i}_{j}=(p^{i}_{j+1}-p^{i}_{j})/\|p^{i}_{j+1}-p^{i}_{j}\|_{2}$ represents the direction of change in strand position, and curvature $c^{i}_{j}=\|o^{i}_{j}-o^{i}_{j-1}\|_{2}$ denote the rate of change in strand orientation. Given a hair description, we select an optimal pre-trained NST as initialization and further optimize it in subsequent stages.
@@
Discussion on Strand-Based Differentiable Rendering. With reasonable hair initialization, the next step is to leverage human-specific 2D diffusion models to further sculpt the hair geometry. However, due to the lack of a stable strand-based differentiable renderer, it is very challenging to optimize hair shape using 2D generative priors like other SDS-based methods. Besides, we expect hair strands to produce mesh-style smooth geometry and texture maps to fully exploit the powerful generative capabilities of 2D diffusion models pre-trained on human mesh data. A feasible alternative is to first differentiably convert the strands into quad meshes [82], similar to NeuralHaircut [64], and then leverage SDS-based mesh optimization frameworks to model hair strands. However, this non-watertight stripe-like mesh easily produces ambiguous normal maps or excessively thin sides (see the zoom image of Fig. 3-(c)), which significantly reduce the optimization stability (see the drifting hair highlighted by the oval dotted box in Fig. 3-(c)). Therefore, a differentiable strand-to-mesh conversion method with superior optimization properties is urgently needed.
Figure 3: The differentiable prismatization algorithm’s (a) gradient backpropagation process, (b) a strand-to-mesh conversion example and (c) advantages over NeuralHaircut [64]. Non-watertight quad meshes can easily produce ambiguous normal maps, which significantly reduce the stability of hair shape modeling (see the drifting hair highlighted by the oval dotted box in (c)).
@@
Differentiable Prismatization (DP). To this end, we propose a novel differentiable prismatization algorithm inspired by the internal structure of 3D hair (i.e., hair fiber is a dielectric cylinder covered with tilted scales and with a pigmented interior [19, 34]). This algorithm can efficiently and differentiably convert strands into watertight prismatic meshes with arbitrary thickness and lateral edges to approximate the cylindrical structure of the hair. This enables smooth backpropagation of gradients from the SDS loss to the 3D strand representation using mesh-based differentiable renderers [22], thereby making it possible to fully exploit 2D generative priors distilled from high-fidelity human mesh data for modeling realistic hair strands. Fig. 3-(a) visualizes the complete gradient backpropagation process.
In specific, given a hair strand $s$, our DP converts it into a watertight prismatic mesh with $K$ lateral edges and radius $R$ through the following five steps: (1) Compute the Initial Normal Vector. (2) Generate $K$ Rotated Normals. (3) Translate to Form Lateral Edges. (4) Construct Lateral Faces. (5) Construct Top and Bottom Faces. An example of this conversion is shown in Fig. 3-(b). Unlike quad meshes [64], our DP effectively avoids ambiguous normals, enabling smooth gradient backpropagation and hair strand optimization, as shown in Fig. 3-(c). Further algorithm details are provided in the Supp. Mat.
With the help of our proposed differentiable prismatization algorithm, we optimize the neural scalp texture $T$ by utilizing rich human-specific 2D generative priors through the following SDS losses (Fig. 2-(b)):
@@
$\nabla_{T}\mathcal{L}^{\text{hn}}_{SDS}$ $=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{\text{hn}}}(n^{\text{h+s}}_{t};y_{\text{h+s}},t)-\epsilon)\frac{\partial n^{\text{h+s}}}{\partial T}\right],$ (3)
@@
$\nabla_{T}\mathcal{L}^{\text{hd}}_{SDS}$ $=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{\text{hd}}}(d^{\text{h+s}}_{t};y_{\text{h+s}},t)-\epsilon)\frac{\partial d^{\text{h+s}}}{\partial T}\right],\vskip-22.76228pt$ (4)
@@
where ${\phi}_{\text{hn}}$ and ${\phi}_{\text{hd}}$ are the human-specific normal-adapted and depth-adapted diffusion models from HumanNorm [17], $n^{\text{h+s}}$ and $d^{\text{h+s}}$ are the rendered normal and depth maps of the head DMTet $\theta_{\text{h}}$ with hair prismatic mesh $M_{p}=DP(G(T))$, respectively. Here, $DP(\cdot)$ denotes the differentiable prismatization operation function, $y_{\text{h+s}}$ is the full text description of the head with hair.
@@
However, due to the excessive flexibility of hair strands, relying solely on the SDS loss leads to unnatural strand orientations and unrealistic hairstyles (Fig. 9). Furthermore, the lack of constrained multi-view images prevents us from leveraging hair-specific features [44] to provide powerful supervision as in previous works [55].
@@
Prior-Driven Losses. To address this issue, we propose two straightforward yet robust prior-driven losses based on observations of the 3D hair strand orientation and curvature to ensure the rationality of the generated hair strands. Specifically, we observe two geometric properties of hair: (1) neighboring strand orientations are highly consistent, and (2) strand curvature is strongly and positively correlated with the curliness of the hairstyle. To validate the properties, we calculate the cosine similarity of adjacent hair strand orientations $CS_{\text{ori}}$ and the average curvature $C_{\text{mean}}$ of 343 hairstyles in the USC-HairSalon dataset [15] using the following equations and visualize their distributions and some examples in Fig. 4:
@@
$C_{\text{mean}}$ $=\frac{1}{N_{s}N_{p}}\sum^{N_{s}}_{i=1}\sum^{N_{p}}_{j=1}c^{i}_{j},\vskip-17.07182pt$ (6)
[…]
Figure 4: Observation of hair geometric features: (1) neighboring strand orientations are highly consistent. (2) strand curvature is strongly and positively related to the haircut curliness.
@@
$\mathcal{L}_{\text{cur}}$ $=\|C_{\text{mean}}-C_{\text{target}}\|_{1},\vskip-14.22636pt$ (8)
where $C_{\text{target}}$ represents the target average curvature set according to the input hair description. These two losses, inspired by 3D hair geometry priors, regularize the hair shape by supervising the consistency of orientations between adjacent strands and the overall hairstyle’s curvature.
Figure 5: Examples of high-fidelity and diverse 3D heads and strand-accurate haircuts generated by our method. The upper visualization includes rendered color and normal maps of the head and hair prismatic meshes. The lower visualization shows the physics-based hair strand rendering result using Blender [10]. For better strand-based visualization, we interpolate generated hair to approximately 10,000 strands and apply a consistent appearance. Please zoom in for detailed views, and refer to the Supp. Mat. for video demonstrations.
Additionally, we introduce a series of losses $\mathcal{L}_{bbox}$, $\mathcal{L}_{face}$, and $\mathcal{L}_{colli}$, which prevent the hair from exceeding the bounding box, obscuring the face, and colliding with the head, respectively. These losses can further improve the rationality of hair geometry. Their details are provided in the Supp. Mat. The final loss function for optimizing hair shape is expressed as follows:
@@
3.5 Hair Texture Generation
Next, we fix the optimized hair strand geometry and model the realistic hair strand texture under the supervision of the SDS loss as follows (Fig. 2-(c)):
$\nabla_{{\psi}_{\text{s}}}\mathcal{L}^{\text{hc}}_{SDS}=\mathbb{E}_{t,\epsilon}({\epsilon}_{{\phi}_{\text{hc}}}(c^{\text{h+s}}_{t};n^{\text{h+s}},y_{\text{h+s}},t)-\epsilon)\frac{\partial c^{\text{h+s}}}{\partial{\psi}_{\text{s}}},\vskip-2.84544pt$ (10)
where ${\phi}_{\text{hc}}$ is the human-specific normal-conditioned diffusion model from HumanNorm [17], $c^{\text{h+s}}$ is the rendered color map of the head texture field ${\psi}_{\text{h}}$ and hair strand texture field ${\psi}_{\text{s}}$. Since the vanilla SDS loss often leads to color oversaturation, we replace it with the following MSDS loss [17] to further enhance the texture’s realism in later iterations of texture optimization:
$\nabla_{{\psi}_{\text{s}}}\mathcal{L}^{\text{hc}}_{MSDS}=\mathbb{E}_{t,\epsilon}(H(c^{\text{h+s}}_{t};n^{\text{h+s}},y_{\text{h+s}},t)-\epsilon)\frac{\partial c^{\text{h+s}}}{\partial{\psi}_{\text{s}}},\vskip-2.84544pt$ (11)
where $H(\cdot)$ denotes the multi-step operation function. Additionally, we propose a strand-aware texture field that models orientation-dependent texture to better generate high-frequency color variations. Please refer to the Supp. Mat. for more their details.
@@ 3.6 Implementation Details
Our method is implemented using PyTorch [46] and ThreeStudio [11]. Experiments are conducted on an Ubuntu server equipped with A6000 GPUs. Starting with a prompt such as “A man with brown hair”, we first generate a bald head using “A bald man”. Then, we generate the hair based on “A man with brown hair”. Further implementation details are available in the Supp. Mat.
@@
Figure 7: Qualitative comparison with HAAR [65]. HAAR, which does not model heads, often produces unreasonable hair-head collisions (highlighted in the black box).
@@ 4 Experiments
Some examples of 3D head models with strand-based hair generated by StrandHead are shown in Fig. 5. Due to space limitations, only the key experimental settings and results are presented here. For more results, evaluations, and details, please refer to our Supp. Mat.
@@ 4.1 Experimental Settings
Baselines. We compare StrandHead with the SOTA methods for head avatar and haircut generation. The baseline methods for head generation include text-to-holistic-head works (HeadArtist [30], HumanNorm [17], and HeadStudio [94]) and text-to-decoupled-head work (TECA [84]). For text-to-hair generation, we consider general text-to-3D methods (MVDream [62], GaussianDreamer [80], LucidDreamer [28], and RichDreamer [48]), text-to-strand method (HAAR [65]), and TECA [84].
@@
Evaluation Metrics. Current text-to-3D methods typically use CLIP-based metrics to assess text-image alignment and output quality. However, research [16, 31] shows that these metrics struggle to capture fine-grained alignment between 3D content and input prompts, as validated by our experiments in the Supp. Mat. To address this limitation, we draw inspiration from Progressive3D [7] and Barbie [66], using BLIP-VQA [24, 23] and BLIP2-VQA [25, 23] for evaluation. Specifically, we convert each prompt into two questions to separately verify head and hair. We then feed the rendered image of the generated 3D content into the VQA model, asking each question in sequence and using the probability of a “yes” response as our evaluation metric. For example, the head prompt “A man with black hair” is transformed into “Is the person in the picture a man?” and “Does the person in the picture have black hair?” The hair prompt “Black hair” becomes “Is the object in the picture black hair?” Finally, we conduct a user study by randomly selecting 10 generated examples and asking 30 volunteers to assess (1) generation quality and (2) text-image alignment, and select the best methods.
@@ 4.2 Comparisons of Head Generation
As illustrated in the upper section of Tab. 2, StrandHead outperforms all comparison methods across every evaluation metric. The qualitative results on the left side of Fig. 6 further highlight the advantages of our method. Compared to other head generation methods, the head avatars generated by StrandHead exhibit not only more refined facial geometry and texture details but also strand-accurate hair with a plausible appearance that integrates seamlessly with physics-based simulation systems. To the best of our knowledge, StrandHead is the first head generation framework that realizes strand-level hair modeling, a capability that holds substantial potential for advancing human-centric 3D AIGC applications in the industry.
@@ 4.3 Comparisons of Hair Generation
Compared with HAAR [65], StrandHead relies on robust 2D/3D human-centric priors rather than large-scale paired datasets to supervise hair generation. This not only allows it to generate hairstyles that are uncommon in the dataset (see slicked-back and side-swept hairstyles in Fig. 7), but also models hair texture or geometry that adapts to specific head shapes (Fig. 8-(b)), thereby avoiding unnatural hair-head collisions (see the black box in Fig. 7).
@@ 4.4 Ablation Study
Figure 8: Analysis of human-specific 2D diffusion models.
Effectiveness of Human-Specific 2D Generative Priors. We demonstrate the importance of human-specific 2D generative priors from two aspects: (1) As shown in Fig. 8-(a), ignoring 2D supervision fails to generate meaningful hair. While a general 2D diffusion model can slightly improve modeling performance, it still falls short of producing high-quality results. Only by incorporating human-specific 2D generative priors can one achieve realistic textures and reasonable shapes for hairstyles. (2) Fig. 8-(b) illustrates generated hair under varying bald head conditions while maintaining a fixed initial hairstyle and hair prompt. The hairstyles created under the supervision of human-specific 2D diffusion models exhibit clear geometric and textural variations that adapt to specific bald heads. This strongly demonstrates the necessity of considering the bald head for hair strand generation and highlights the advantages of StrandHead over HAAR [65].
Figure 9: Ablation study on (a) orientation consistency loss $\mathcal{L}_{\text{ori}}$ and (b) curvature regularization loss $\mathcal{L}_{\text{cur}}$.
Effectiveness of Prior-Driven Losses. Fig. 9-(a) demonstrates that incorporating $\mathcal{L}_{\text{ori}}$ provides strong guidance for generating natural hair by supervising the consistency of orientations between adjacent strands. As shown in Fig. 9-(b), our $\mathcal{L}_{\text{cur}}$ preserves the desired curliness by constraining the overall hairstyle’s curvature. In summary, these two losses, inspired by 3D hair geometry priors, effectively enhance the realism and rationality of the generated haircut.
Figure 10: Failure cases.